### v1

@@ Abstract
The problem of using structured methods to represent knowledge is well-known in conceptual modeling and has been studied for many years. It has been proven that adopting modeling patterns represents an effective structural method. Patterns are, indeed, generalizable recurrent structures that can be exploited as solutions to design problems. They aid in understanding and improving the process of creating models. The undeniable value of using patterns in conceptual modeling was demonstrated in several experimental studies. However, discovering patterns in conceptual models is widely recognized as a highly complex task and a systematic solution to pattern identification is currently lacking. In this paper, we propose a general approach to the problem of discovering frequent structures, as they occur in conceptual modeling languages. As proof of concept for our scientific contribution, we provide an implementation of the approach, by focusing on UML class diagrams, in particular OntoUML models. This implementation comprises an exploratory tool, which, through the combination of frequent subgraph mining algorithm and graph manipulation techniques, can process multiple conceptual models and discover recurrent structures according to multiple criteria. The primary objective is to offer a support facility for language engineers. This can be employed to leverage both good and bad modeling practices, to evolve and maintain the conceptual modeling language, and to promote the reuse of encoded experience in designing better models with the given language.
@@
Recently, approaches to facilitate this complex discovery process produced valuable results [8, 9]. An empirical study [10] demonstrated that the Frequent Subgraph Mining algorithm [11] can be used to automatically discover certain kinds of patterns in OWL vocabularies. Works like [8] showed how the implementation of Tree Mining [12] techniques aids the detection of specific types of patterns in BPMN. Still, a general solution to this problem has not been devised.
In this paper, we present an approach for facilitating the empirical discovery of recurrent modeling structures. We implemented this approach in a tool that acts as an exploratory facility, which allows users to discover and analyze patterns interactively. The tool relies on the application of gSpan [13], a well-known frequent subgraph mining algorithm. Through a command-line interface, it allows users to:
@@
prepare the data to be mined by focusing on specific information and filtering out concepts that might be not relevant for that type of analysis,
@@
select certain features that the output patterns should have (e.g., number of nodes, frequency, or dissimilarities from known patterns),
@@
visualize the output patterns and query the input models to verify the domain occurrences of that pattern, and
@@
cluster the output patterns to simplify the user’s final assessment.
The main scope of our solution is to offer a support facility for language engineers that aims at exploiting bad/good practices to evolve and maintain the conceptual modeling language, but also to favor the reuse of encoded experience in designing better models with the given language.
We evaluate our approach using a dataset of UML Class Diagrams, in particular OntoUML models. We adopted this language for several reasons. Primarily, OntoUML is a language based on pre-defined patterns, which are frequently used across different models to address common modeling issues. This can allow us to evaluate our approach to finding useful structures that are already known by the designers of the language itself. Second, for our evaluations, we can access a curated dataset of OntoUML models [14], encoded in a uniform format. Third, for our evaluations, we can obtain feedback directly from the authors of the language, enabling us to gather valuable information about the practical utility of the approach. Finally, OntoUML has a variety of constructs and features that allow us to assume that if our approach works for this language, it also works for widely used standard languages, such as UML or BPMN.
The structure of this paper is as follows: Section 2 explores the concept of “pattern” in conceptual modeling, focusing particularly on “recurrent modeling structures”. This section also provides a brief overview of the primary technique employed for automating the mining process. Section 3 enumerates the requirements that guided the design of our methodology. In Section 4, we introduce our framework, elaborating on its workflow and individual components. Section 5 details the implementation of our approach. In Section 6 and Section 7, we discuss the experiments and demonstrations conducted to validate our solution. In Section 8, we analyze insights gained from the evaluation and demonstration processes. Section 9 situates our work within the context of existing literature. Finally, Section 10 offers reflections on our findings.
@@ 2.1 Patterns in Conceptual Modeling
Currently, our proposal primarily focuses on patterns in conceptual models. These model fragments represent recurrent structures formed by modeling constructs from a specific conceptual modeling language. Keeping track of these patterns allows one to understand how a language is applied, thus offering powerful means to language designers for maintaining and evolving the language itself. Using recurrent structures, one can, e.g., discover the most common combinations of language constructs, identify language dialects for specific application domains, verify possible restrictions in using constructs or their combinations, and determine the frequent subversions of the language.
The notion of systematic language subversion [15] refers to an ungrammatical use of a language’s constructs that becomes recurrent in a language community signalling a design limitation of that language. It is closely related to coding traditions [16] that cover coding policies, notational guidelines, naming conventions, implementation patterns, programming idioms, etc.
Figure 1 shows examples of common usages of the generalization relationship in UML class diagrams. These structures represent the ways by which modelers represent intentional sub-typing (inheritance) between classes (see .a, .b and .c types), and can be taken as useful insights for understanding the language application. For instance, given a set of models about a certain domain of information, it is possible to discover that the multiple inheritance (.c) is used to capture class compositionally [17]. This observation might trigger language designers to suggest class composition relations when modelers use two or more generalization relations for a class.
@@ 2.2 Frequent Subgraph Mining
FSM presents two distinct problem formulations: .i graph transaction-based FSM, and .ii single graph-based FSM. In graph transaction-based FSM, the input dataset comprises a collection of medium-sized graphs termed “transactions” [19]. On the other hand, single graph-based FSM uses a single, significantly large graph as input. In this setting, a subgraph, denoted as g, is deemed frequent if its occurrence count surpasses a predefined threshold value, commonly referred to as the support threshold. The support calculation varies according to the problem formulation. In a transactional dataset, the support of a graph is the ratio of the number of transactions to which this graph occurs to the total number of transactions. In a single large graph, the support of a graph is the number of its occurrences in this dataset.
Formally, given a graph , where is a set of (possibly labeled) vertices and is a set of (possibly labeled) edges, then a graph is a sub-graph of G if and only if its vertices and edges are a subset of the vertices and edges of graph G. A subgraph is considered frequent when its support equals or exceeds the user-defined minimum support threshold.
Let us consider a graph database represented in Figure 3. Suppose that we want to discover all the (connected) subgraphs with at least two nodes that occur in at least three graphs. By adopting minimum support 3 as a parameter and applying an FSM algorithm, we can obtain the set of all subgraphs appearing in at least three graphs, namely .
In practical scenarios, different tools or algorithms can identify frequent subgraphs. In this paper, we adopt gSpan [13], a popular and important state-of-the-art solution. This algorithm is often used for discovering frequent subgraphs in large graph databases, and has variants that work on conceptual graphs [20], typed linked data [21], chemical compound data [22], large disk-based databases [23], data with differential privacy [24], etc. It uses an approach based on a minimum DFS code [25] that allows to efficiently generate candidate subgraphs and prune infrequent ones. The algorithm iteratively discovers all frequent subgraphs containing an edge and then expands the search to include larger subgraphs. The database is shrunk as the search continues, and only graphs containing the current subgraph are considered. Studies show that gSpan can mine large frequent subgraphs with lower minimum supports and high performance [11].
@@ 3 Requirements
Our design for the proposed approach is based on an initial problem identification activity. During this phase, we gathered feedback from five potential users of our approach: expert language creators, who have been involved in the development of different conceptual modeling languages and who have been working on the identification of conceptual modeling patterns. We performed open-ended interviews and the main open questions we asked were:
@@
Interestingness: The approach should facilitate the discovery of subjectively interesting patterns, namely recurrent structures that can be considered interesting according to the user’s interpretation. Here, the notion of “subjectively interesting” is inspired by the work from Silberschatz and Tuzhilin [26], where a pattern is ranked as interesting by a user mainly because .i it is considered exploitable for the modeling activities, or .ii it contradicts some user’s expectations.
@@
Comprehension: The approach should assist engineers in the process of assessing and analyzing the output patterns. This should be feasible by accounting for multiple types of frequency, e.g., relative frequency, as the number of occurrences of a pattern in each model; total frequency, as the overall occurrences for a given pattern; frequency across models, as the number of models in which a pattern occurs. Moreover, this should be supported by a human-readable visualization format of the output patterns.
@@ 4.1 Preparing the Input Data
The “Preparation” phase focuses on transforming the input conceptual models in a format that is processable by the subsequent steps. In this phase, we have two main tasks.
@@
First, the Importing (0) task consists of taking a set of conceptual models encoded in a language (e.g., UML or BPMN) and transforming each model into a Labeled Property Graph (LPG) .111Note that this task is language-dependent and requires an ad hoc transformation for each source conceptual modeling language. At first glance, this might seem straightforward since conceptual models are essentially graphs. That, however, is not always the case. Consider, for instance, the transformation of UML class diagrams into graphs. The simple solution is transforming classes into nodes and generalizations and associations into edges. Still, if we want to convert, generalization sets, association classes, generalizations between associations, cardinalities, and several other constructs, this solution no longer works. Moreover, a challenge for the available frequent subgraph mining approaches is accounting for complex (semantically-rich) graph data [11, 29, 30]. There are several algorithms each of which admits graphs with different characteristics (e.g., labeled vs. unlabeled or directed vs. undirected). In our case, we have to account for highly expressing modeling languages, such as OntoUML, which represent limiting cases that challenge the functionality of the available mining solutions. Therefore, we opted for full compatibility with all kinds of algorithms, solving the problem at the source and generating a format that is always processable.
@@
Data: Set of of Conceptual Models Result: Set of Labeled Property Graphs (LPG) 1 for every conceptual model , where do 2 map each class into a node ; 3 map each association into a anode ; 4 map each generalization into a node ; 5 map association cardinalities into a node ; 6 for each generalization node do 7 Connect to the parent class-node with an edge “general”; 8 Connect to the child class-node with an edge “specific”; 9 end for 10 for each association node do 11 Connect to the source class-node with an edge “source”; 12 Connect to the target class-node with an edge “target”; 13 for each cardinalities node do 14 Connect to the corresponding with an edge “cardinalities”; 15 end for 16 17 end for 18 for each class do 19 create the corresponding label(s) for node ; 20 end for 21 for each association do 22 create the corresponding label(s) for node ; 23 24 end for 25 add a graph to ; 26 27 end for 28
Algorithm 1 offers a partial view of how the importing step works. The main rationale is then to reify everything that is in the input conceptual model. For instance, associations, cardinalities, and classes are all mapped into graph nodes with multiple labels. Figure 2 provides an example of transformation occurring via importing. .a represents the input conceptual model, encoded into OntoUML, .b represents the output labeled property graph. The edge labels “source-target” and “specific-general” are adopted to keep track of the directionality of generalizations and associations, respectively.
@@
4.2 Discovering Patterns Across and Within Models
The “Discovery” phase is the core of the workflow and is aimed at generating the candidate patterns in a format that can be then processed and made accessible to the user for the final assessment. This phase, in turn, is composed of two main tasks.
@@ 4.2 Discovering Patterns Across and Within Models
the minimum pattern size for the output patterns (e.g., filter out patterns that have less than five nodes),
@@
the known patterns to be excluded from the final output (e.g., the user may want to filter out subgraphs that are highly expected to focus on unexpected patterns).
@@
4.3 Assessing the Output Patterns
@@
The Clustering (4) task groups output patterns—specifically, the recurrent or frequent structures from the mining task—based on their similarities.333Notice that in the approach we are proposing, to assess whether one conceptual model sub-structure is close to another, we take inspiration from conceptual model similarity techniques [31, 32], where the main task is to find similar models given a reference model input, to categorizing models according to their characteristics. This should facilitate the user’s process of scraping and consulting the output. Often, the number of outputs produced can surpass hundreds, and numerous patterns can have similar characteristics. For example, some structures might differ by just one label (in one structure we might have three nodes where one is labeled as Kind and the other two specialize it as Subkind, while in another structure we might have three nodes where one is Kind, and the other two specialize it as Subkind and Role, respectively. In this sense, a cluster of patterns can be considered by the user as a set of variations, which sometimes may refer also to the same structural pattern (e.g., where a kind is specialized by two or three Subkinds). To enable clustering, in our process, we extract key characteristics from the graphs that represent each pattern. At present, the steps addressed to enable the clustering step are as from Algorithm 2 below.
Currently, we adopt a single feature extraction method to calculate pattern similarity. However, this does not prevent adopting different feature extraction or embedding techniques. This is valid also for the approach used to assess similarity, which now is limited to a cosine similarity [33], but in the future, it can be extended to other measures.
@@
The significance of this visual representation lies in its ability to empower users in several ways. For example, by translating the patterns from the FSM algorithm into multi-directed graphs, users can better grasp complex relationships and hierarchies within the data. This enables deeper insights and understanding. Figure 3 exemplifies the transformation power of visualization. On the left-hand side, you can observe a recurrent structure. On the right, you are presented with one of its occurrences, generated through the deepening task. This visual juxtaposition illustrates how .a can be transformed into a visually accessible and actionable representation, empowering users to extract valuable knowledge from their data with ease.

### v2

@@ Abstract
The challenge of using structured methods to represent knowledge is a well-documented issue in conceptual modeling and has been the focus of extensive research. It is widely recognized that adopting modeling patterns offers an effective structural approach for designing conceptual models. Patterns, in this context, refer to generalizable, recurring structures that provide solutions to common design problems. They significantly enhance both the understanding and improvement of the modeling process. Numerous experimental studies have demonstrated the undeniable value of using patterns in conceptual modeling. Despite this, the task of identifying patterns in conceptual models remains highly complex, and there is currently no systematic method for pattern discovery. To address this gap, this paper proposes a general approach for discovering frequent structures in conceptual modeling languages as a means to support pattern identification. Specifically, we focus on uncovering recurring structures that reflect the usage patterns of a given conceptual modeling language. As proof of concept, we implement our approach by focusing on two widely-used conceptual modeling languages. This implementation includes an exploratory tool that integrates a frequent subgraph mining algorithm with graph manipulation techniques. The tool processes multiple conceptual models and identifies recurrent structures based on various criteria. We validate the tool using two state-of-the-art curated datasets: one consisting of models encoded in OntoUML and the other in ArchiMate. The primary objective of our approach is to provide a support tool for language engineers. This tool can be used to identify both effective and ineffective modeling practices, enabling the refinement and evolution of conceptual modeling languages. Furthermore, it facilitates the reuse of accumulated expertise, ultimately supporting the creation of higher-quality models in a given language.
@@
The analysis of conceptual patterns and the discovery of recurring structures or useful patterns has emerged as a distinct research path that has seen significant growth in recent years. This research, as evidenced by seminal works like [8, 6, 7, 9], focuses on analyzing reference patterns—used as case studies—to infer modeling strategies for specific problems. More recent work has explored the application of automated techniques to support heuristic tasks traditionally performed manually [10, 11, 12, 13].
In this article, we join the ongoing research efforts to develop automated support for facilitating the empirical discovery of recurring modeling structures. Our contribution is an approach that integrates several state-of-the-art techniques, designed as an exploratory tool to enable users to interactively discover and analyze frequent structures. The implementation of this approach is based on gSpan [14], a widely recognized frequent subgraph mining algorithm. Through a command-line interface, this tool enables users to:
@@
prepare the data to be mined by focusing on specific information and filtering out concepts that might not be relevant for that type of analysis,
@@
select certain features that the output frequent structures should have (e.g., number of nodes, frequency, or dissimilarities from known patterns),
@@
visualize the output frequent structures and query the input models to verify the domain occurrences of that structure, and
@@
cluster the output structures to simplify the user’s final assessment.
The main scope of our solution is to offer a support facility for language engineers that aims at exploiting bad/good practices to evolve and maintain the conceptual modeling language but also to favor the reuse of encoded experience in designing better models with the given language.
We evaluate our approach using two large and curated state-of-the-art datasets of frequently used conceptual modeling languages, in particular, the OntoUML dataset [15] and the ArchiMate dataset [16]. We adopted datasets encoded in OntoUML and ArchiMate languages for several reasons. Primarily, these languages are based on pre-defined patterns, which are frequently used across different models to address common modeling issues. This allows us to evaluate our approach to finding useful structures that are already known by the designers of each language. Second, for our evaluations, we can access curated high-quality datasets, encoded in a uniform format, thus improving our findings’ comprehensibility and reliability. Third, for our evaluations and experiment design, we can obtain feedback directly from the authors of the languages,111In particular, we leveraged this feedback in the OntoUML scenario. showing how this may help enabling us to gather valuable information about the practical utility of the approach. Finally, OntoUML and ArchiMate have a variety of constructs and features that allow us to assume that if our approach works for these languages, it also works for widely used standard languages, such as UML or BPMN.
The outline of this paper is as follows: Section 2 explores the concept of “pattern” in conceptual modeling, focusing particularly on “recurrent modeling structures”. This section also provides a brief overview of the primary technique employed for automating the mining process. Section 3 enumerates the requirements that guided the design of our methodology. In Section 4, we introduce our framework, elaborating on its workflow and individual components. Section 5 details the implementation of our approach. In Section 6 and Section 7, we discuss the experiments and demonstrations conducted to validate our solution. In Section 8, we analyze insights gained from the evaluation and demonstration processes. Section 9 situates our work within the context of existing literature. Finally, Section 10 offers reflections on our findings.
@@ 2.1 Patterns in Conceptual Modeling
Currently, our proposal primarily focuses on patterns in conceptual models. These model fragments represent recurrent structures formed by modeling constructs from a specific conceptual modeling language. Keeping track of these patterns allows one to understand how a language is applied, thus offering powerful means to language designers for maintaining and evolving the language itself. Using recurrent structures, one can, e.g., discover the most common combinations of language constructs, identify language dialects for specific application domains, verify possible restrictions in using constructs or their combinations, and determine the frequent subversions of the language. The notion of systematic language subversion [17] refers to an ungrammatical use of a language’s constructs that becomes recurrent in a language community signalling a design limitation of that language. It is closely related to coding traditions [18] that cover coding policies, notational guidelines, naming conventions, implementation patterns, programming idioms, etc.
Figure 1 shows examples of common usages of the generalization relationship in UML class diagrams. These structures represent the ways by which modelers represent intentional sub-typing (inheritance) between classes (see .a, .b and .c types), and can be taken as useful insights for understanding the language application. For instance, given a set of models about a certain domain, it is possible to discover that the multiple inheritance (.c) is used to capture class compositionally [19]. This observation might trigger language designers to suggest class composition relations when modelers use two or more generalization relations for a class.
@@ 2.2 Frequent Subgraph Mining
FSM presents two distinct problem formulations: .i graph transaction-based FSM, and .ii single graph-based FSM. In graph transaction-based FSM, the input dataset comprises a collection of medium-sized graphs termed “transactions” [22]. On the other hand, single graph-based FSM uses a single, significantly large graph as input. In this setting, a subgraph, denoted as g, is deemed frequent if its occurrence count surpasses a predefined threshold value, commonly referred to as the support threshold. The support calculation varies according to the problem formulation. In a transactional dataset, the support of a graph is the ratio of the number of transactions to which this graph occurs to the total number of transactions. In a single large graph, the support of a graph is the number of its occurrences in this graph.
Formally, given a graph , where is a set of (possibly labeled) vertices and is a set of (possibly labeled) edges, then a graph is a sub-graph of G if and only if its vertices and edges are a subset of the vertices and edges of graph G, and the vertex subset include all endpoints of the edge subset. A subgraph is considered frequent when its support equals or exceeds the user-defined minimum support threshold.
Let us consider the three graphs represented in Figure 3. Suppose that we want to discover all the (connected) subgraphs with at least two nodes that occur in at least three graphs. By adopting minimum support 3 as a parameter and applying an FSM algorithm, we can obtain the set of all subgraphs appearing in at least three graphs, namely .
In practical scenarios, different tools or algorithms can identify frequent subgraphs. In this paper, we adopt gSpan [14], a popular state-of-the-art solution. This algorithm is often used for discovering frequent subgraphs in large graph databases and has variants that work on conceptual graphs [23], typed linked data [24], chemical compound data [25], large disk-based databases [26], data with differential privacy [27], etc. It uses an approach based on a minimum DFS code [28] that allows the efficient generation of candidate subgraphs and pruning of infrequent ones. The algorithm iteratively discovers all frequent subgraphs containing an edge and then expands the search to include larger subgraphs. The database is shrunk as the search continues, and only graphs containing the current subgraph are considered. Studies show that gSpan can mine large frequent subgraphs with lower frequency and high performance [21].
@@ 3 Requirements
Our design for the proposed approach is based on an initial problem identification activity. During this phase, we gathered feedback from five potential users of our approach: expert language engineers, who have been involved in the development of different conceptual modeling languages and who have been working on the identification of conceptual modeling patterns. We performed open-ended interviews and the main open questions we asked were:
@@
Interestingness: The approach should facilitate the discovery of subjectively interesting patterns, namely recurrent structures that can be considered interesting according to the user’s interpretation. Here, the notion of “subjectively interesting” is inspired by the work from Silberschatz and Tuzhilin [29], where a pattern is ranked as interesting by a user mainly because .i it is considered exploitable for the modeling activities, or .ii it contradicts some users’ expectations.
@@
Comprehension: The approach should assist engineers in the process of assessing and analyzing the output structures. This should be feasible by accounting for multiple types of frequency, e.g., relative frequency, as the number of occurrences of a structure in each model; total frequency, as the overall occurrences for a given pattern; frequency across models, as the number of models in which a pattern occurs. Moreover, this should be supported by a human-readable visualization format of the output structures.
@@ 4.1 Preparing the Input Data
The “Preparation” phase focuses on transforming the input conceptual models in a format that is processable by the subsequent steps. In this phase, we have two main tasks. First, the Importing (0) task consists of taking a set of conceptual models encoded in a language (e.g., UML or BPMN) and transforming each model into a Labeled Property Graph (LPG) . Such a task is denoted by a white box in Figure 4 because it is language-dependent, namely it requires an ad hoc transformation for each source conceptual modeling language taken into consideration. At first glance, this might seem straightforward since conceptual models are essentially graphs. That, however, is not always the case. Consider, for instance, the transformation of UML class diagrams into graphs. The simple solution is transforming classes into nodes and generalizations and associations into edges. Still, if we want to convert, generalization sets, association classes, generalizations between associations, cardinalities, and several other constructs, this solution no longer works. Moreover, a challenge for the available frequent subgraph mining approaches is accounting for complex (semantically-rich) graph data [21, 32, 33]. There are several algorithms each of which admits graphs with different characteristics (e.g., labeled vs. unlabeled or directed vs. undirected).
In the conceptual modeling context, we encounter construct-rich languages that challenge the capabilities of existing FSM solutions. For instance, some languages support multiple labels for classes and relations, multiple relationships between the same classes, and even classes of relations. This requires the mining algorithm to potentially handle various types of graphs, such as multi-labeled, directed, or multi-graphs. To the best of our knowledge, no existing algorithm fully accommodates all these graph types. Therefore, we ensured full compatibility with all algorithms by addressing the problem at its source. The importing process generates an LPG that captures all the information from conceptual models encoded in any conceptual modeling language, ensuring the output format is always suitable for the mining algorithm.
@@
Data: Set of Conceptual Models Result: Set of Labeled Property Graphs (LPG) 1 for each conceptual model do 2 for each concept in do 3 map to a node ; 4 if represents a relation then 5 Connect to the source node with an edge labeled “source”; 6 Connect to the target node with an edge labeled “target”; 7 8 end if 9 for each property of do 10 assign a label to node ; 11 12 end for 13 14 end for 15 create an undirected labeled graph using the set of nodes and edges; 16 add to ; 17 18 end for 19
Algorithm 1 presents a simplified overview of the importing process. The core idea is to reify every element in the input conceptual model. For example, associations, cardinalities, and classes are all mapped into graph nodes with multiple labels. Figure 5 illustrates this transformation process, applied to three models encoded in different modeling languages: (a) OntoUML, (b) ArchiMate, and (c) BPMN. On the right, the resulting LBGs are shown. The edge labels “source” and “target” are used to preserve the directionality of the relationships.
@@
4.2 Discovering Structures Across and Within Models
The “Discovery” phase is the workflow’s core and aims to generate the candidate patterns in a format that can then be processed and made accessible to the user for the final assessment. This phase, in turn, is composed of two main tasks.
@@ 4.2 Discovering Structures Across and Within Models
the minimum structure size for the output patterns (e.g., filter out patterns that have less than five nodes),
@@
the known patterns to be excluded from the final output (e.g., the user can provide graphs representing known patterns as input. These will be excluded from the final output to maximize the discovery of results that contain new information).
@@
4.3 Assessing the Output Frequent Structures
@@
The Clustering (4) task groups output structures—specifically, the recurrent or frequent structures from the mining task—based on their similarities.333Notice that in the approach we are proposing, to assess whether one conceptual model sub-structure is close to another, we take inspiration from conceptual model similarity techniques [34, 35, 36], where the main task is to find similar models given a reference model input, to categorizing models according to their characteristics. This should facilitate the user’s process of scraping and consulting the output. Often, the number of outputs produced can surpass hundreds, and numerous structures can have similar characteristics. For example, some structures might differ by just one label (in one structure we might have three nodes where one is labeled as Kind and the other two specialize it as Subkind, while in another structure we might have three nodes where one is Kind, and the other two specialize it as Subkind and Role, respectively. In this sense, a cluster of patterns can be considered by the user as a set of variations, which sometimes may refer also to the same structural pattern (e.g., where a kind is specialized by two or three Subkinds). To enable clustering, in our process, we extract key characteristics from the graphs that represent each pattern. At present, the steps addressed to enable the clustering step are as from Algorithm 2.
Currently, we adopt a single feature extraction method to calculate structure similarity. However, this does not prevent adopting different feature extraction or embedding techniques. This is valid also for the approach used to assess similarity, which now is limited to a cosine similarity [37], but in the future, it can be extended to other measures.
@@
The significance of this visual representation lies in its ability to empower users in several ways. For example, by translating the outputs from the FSM algorithm into multi-directed graphs, users can better grasp complex relationships and hierarchies within the data. This enables deeper insights and understanding. Figure 6 exemplifies the transformation power of visualization. On the left-hand side, you can observe a recurrent structure. On the right, you are presented with one of its occurrences, generated through the deepening task. This visual juxtaposition illustrates how the raw outputs of the mining and deepening tasks can be transformed into a visually accessible and actionable representation, empowering users to easily extract valuable knowledge from their data.